OpenStudy (anonymous):

How do you find one sided limits without using a table of value? I.e: Limit as x->3 on its left hand side (3-) of x^2 l x-3 l / x -3

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

Now one sided limits are apllied when the function values behave differently when approaching the point from the left versus the right hand side

OpenStudy (anonymous):

What is the very clear factor there that behaves differently ?

OpenStudy (anonymous):

Meaning around the poiny x =3 - just to the left vs. just to the right ...

OpenStudy (anonymous):

@MarcLeclair ?

OpenStudy (anonymous):

but how do you end up solving it without using a table of value, I can manage to find the limits of the right hand side using a table of value, I can't work it out algebraically. Unless there is no way other than error and trial

OpenStudy (anonymous):

You very easily will understand that u CAN algebraically - just answer my guiding question

OpenStudy (anonymous):

What factor changes behaviour around x=3

OpenStudy (anonymous):

The Y values? or is it that the function isn't continuous like piece wise functions or step functions?

OpenStudy (anonymous):

It is continuous but it ALSO is kind-of piece-wise defined yes - two definition around 3

OpenStudy (anonymous):

So the left hand side limit is different than the right hand side, lets just say at random numbers the right hand side is f(x) = 3 and left hand side is f(x)=10 (random values that i picked)

OpenStudy (anonymous):

Marc please concentrate on my guiding question: WHAT ALGEBRAIC FACTOR CHANGES BEHAVIOR AROUND 3 ?

OpenStudy (anonymous):

I can't tell you, any hint? I am trying to think but i really can't figure it out.

OpenStudy (anonymous):

|x| = \[|t| = \left\{ t..... IF... t> 0 \right\}\] \[|t| = -t ......IF... t<0\]

OpenStudy (anonymous):

|x-3| = x-3 for x >3 AAand it changes to |x-3| = -(x-3) for x<3

OpenStudy (anonymous):

SO it follows that when you divide by (x-3) factor then u get

OpenStudy (anonymous):

Oh! I am so sorry I didn't find the question right away. So the limit changes at 3 because the absolute value makes it either + or -. Therefore, when removing the absolute value bracket , you are left with 2 options. either - (x -3 ) or (x +3 )

OpenStudy (anonymous):

\[\frac{ (x-3) }{ x-3 } = 1\]

OpenStudy (anonymous):

1 is on the right of course

OpenStudy (anonymous):

so then you have to do the opposite so - (x -3) to get on the left hand side!

OpenStudy (anonymous):

Aaand \[\frac{ -(x-3) }{ (x-3) } = -1\] on the left hand side

OpenStudy (anonymous):

therefore it will give you - 1 and then you will end up with -9

OpenStudy (anonymous):

thank you very much for your patience!

OpenStudy (anonymous):

As they say in Kentucky : Yep, ya right on it

OpenStudy (anonymous):

can i ask you another question on here really quickly?

OpenStudy (anonymous):

click the blue ribbon "best response"

OpenStudy (anonymous):

Ask

OpenStudy (anonymous):

What if the question was without an absolute so lets say lim x-> -3 on the left hand side of x-3 / x+ 3. I can't factor anything and the denominator makes me end up with 0. So the answer would be infinity?

OpenStudy (anonymous):

NOO! Algebra IS allowed - algebraically your FRACTION is ODENTICALLY 1 everywhere except 3 where it is undefined (because of 0). But limit does not care about the single point. Limit of a constant 1 = 1

OpenStudy (anonymous):

Thank you! :)